OpenStudy (anonymous):

O and O are similar O-rings. The inner radius of O is 5 ft and the inner radius of O is 15 ft a) if the circumference of the outer circle of O is 14π ft what is the circumference of the outer circle of O? b) suppose that is takes 1.5 gallons of paint to paint the O ring O. If the paint is used at the same rate. how much paint is needed to paint the O-ring O?

OpenStudy (anonymous):

This problem is confusing. Are the two different rings really given the same variable name or was that a problem with copying/pasting the question?

OpenStudy (anonymous):

same variable name its dealing with similarity

OpenStudy (anonymous):

C1/ c2 = 2 * 15 pi / 2 pi 5 = 3 -> C1 = 3 * 14 pi = 42 pi